# AdamX: An Adam improvement algorithm based on a novel exponential decay mechanism for the second-order moment estimate

### v2

@@ Abstract
Since the 21st century, artificial intelligence has been leading a new round of industrial revolution. Under the training framework, the optimization algorithm aims to stably converge high-dimensional optimization to local and even global minima. Entering the era of large language models, although the scale of model parameters and data has increased, Adam remains the mainstream optimization algorithm. However, compared with stochastic gradient descent (SGD) based optimization algorithms, Adam is more likely to converge to non-flat minima. To address this issue, the AdamX algorithm is proposed. Its core innovation lies in the proposition of a novel type of second-order moment estimation exponential decay rate, which gradually weakens the learning step correction strength as training progresses, and degrades to momentum SGD in the stable training period, thereby improving the stability of training in the stable period and possibly enhancing generalization ability. Experimental results show that our second-order moment estimation exponential decay rate is better than the current second-order moment estimation exponential decay rate, and AdamX can stably outperform Adam and its variants in terms of performance. Our code is open-sourced at https://github.com/mengzhu0308/AdamX.
@@
Optimization algorithms , Adam , SGD , AdamX , Exponential decay rate of second-order moment estimate.
@@ 1 Introduction
With the advent of the era of large language models (LLM), the currently mainstream optimization algorithm remains Adam. Therefore, it still holds research value to make some minor modifications to Adam. The main criticisms of Adam are twofold: (1) it requires additional sets of parameters to cache the first-order and second-order moments. (2) It tends to converge to sharp minima, which have poorer generalization capabilities [13, 14]. In response to the issue of Adam’s slightly inferior generalization, this paper proposes the AdamX algorithm (see Algorithm 1), whose core innovation lies in the proposition of a novel exponential decay rate for the second-order moment estimation, which controls the gradual weakening of the learning step correction strength as training progresses, thereby degrading to momentum SGD in the stable training period.
The remainder of this paper is organized as follows. Section 2 reviews the existing research on optimization algorithms in deep learning and introduces the motivation of this study. Section 3 provides a detailed analysis of the problem and proposes the corresponding solution. Sections 4, 5, and 6 jointly verify the superiority of AdamX through experiments. The final section summarizes the content of the entire paper.
@@ 3 Proposed algorithm
This section first analyzes the impact of $\beta_{2}$ on the variance of the second-order moment estimation in Adam, semi-qualitatively and semi-quantitatively proving that when $t\to\infty,\beta_{2}\to 1$, Adam can degrade into momentum SGD, thereby reducing the “jitter” phenomenon in the stable training period. Then, to address the contradiction in Adam that $\beta_{2}$ cannot be too large or too small, a novel exponential decay rate for the second-order moment estimation is proposed, along with the corresponding AdamX optimization algorithm. Finally, AdamX and Adam are compared once again.
@@ 3.1 Impact of $\beta_{2}$ on the variance of the second-order moment estimation in Adam
3.2 Novel exponential decay rate for the second-order moment estimation and AdamX
@@
and correspondingly proposes the AdamX algorithm, as shown in Algorithm 1. To align with Adam’s default settings: $\beta_{1}=0.9$, $\beta_{2}=0.999$, AdamX is set to: $\beta_{1}=0.9$, $\beta_{2}=0.99$. This allows hyperparameters tuned for Adam to be better transferred to AdamX. Figure 1 visualizes $\overset{\frown}{\beta}_{1,t}$ and $\overset{\frown}{\beta}_{2,t}$ in AdamX.
@@ 13    $\mathcal{L}_{0}\leftarrow\mathcal{L}_{t}$;
Algorithm 1 AdamX algorithm
Figure 1: Functional curves of $\overset{\frown}{\beta}_{1,t}$ and $\overset{\frown}{\beta}_{2,t}$ versus iteration number $t$ in AdamX
3.3 AdamX v.s. Adam
@@
Comparing Equations (12) and (14), the core difference is that when $t\to\infty$, AdamX’s $\overset{\frown}{\beta}_{2,t}\to 1$, thus AdamX degenerates to momentum SGD, guaranteeing more stability in the later training stage. Figure 2 further visualizes the comparison between AdamX’s $\overset{\frown}{\beta}_{2,t}$ and Adam’s $\overset{\frown}{\beta}_{2,t}$ for different $\beta_{2}$ values.
Figure 2: Comparison between AdamX’s $\overset{\frown}{\beta}_{2,t}$ and Adam’s $\overset{\frown}{\beta}_{2,t}$
@@ 5 Comparison results and analysis of optimization algorithms of the same type
Firstly, the superiority of AdamX in performance was verified by comparing it with optimization algorithms of the same type (Adam [2] and AdaX [22]) (These optimization algorithms can use the same training configuration). Experiments were conducted on the CIFAR-100 [24], PASCAL VOC [25], and Semantic Boundaries [26] datasets, with EfficientNetV2 [27] and SwinV2 [28] selected as the benchmark models for image classification, YOLOv7-tiny [29] as the benchmark model for object detection, and U-Net [30] as the benchmark model for semantic segmentation.
@@ 5.1 Image classification experimental results and analysis
Lowest error rate comparison. Table 3 shows the comparative experimental results on the benchmark model EfficientNetV2-M. The total number of training epochs was set to $80$, the batch size was set to $64$, the peak learning rate was set to $3\times 10^{-3}$, the minimum learning rate was set to $3\times 10^{-5}$, and the training $t_{1}$ was set to $37,488$. As seen in Table 3, our AdamX achieved the lowest top-1 error rate of $34.27\%$ and the lowest top-5 error rate of $11.40\%$ on EfficientNetV2-M. Table 4 shows the comparative experimental results on the benchmark model SwinV2-S. The total number of training epochs was set to $80$, the batch size was set to $64$, and the fixed learning rate was set to $2.2\times 10^{-5}$. As seen in Table 4, our AdamX also achieved the lowest top-1 error rate of $39.42\%$ and the lowest top-5 error rate of $14.68\%$ on SwinV2-S.
@@ 5.2 Object detection experimental results and analysis
Highest mAP comparison. Table 5 shows the comparative experimental results on the VOC2007 test benchmark dataset. The total number of training epochs was set to $80$, the batch size was set to $32$, the peak learning rate was set to $3\times 10^{-3}$, the minimum learning rate was set to $3\times 10^{-5}$, and the training $t_{1}$ was set to $25,680$. As seen in Table 5, our AdamX achieved the highest mAP@0.5 of $52.18\%$ and the highest mAP@0.75 of $28.21\%$.
@@ 5.3 Semantic segmentation experimental results and analysis
Highest mIoU comparison. Table 6 shows the comparative experimental results on the Semantic Boundaries benchmark dataset. The total number of training epochs was set to $80$, the batch size was set to $32$, the peak learning rate was set to $1\times 10^{-3}$, the minimum learning rate was set to $1\times 10^{-5}$, and the training $t_{1}$ was set to $12,720$. As can be seen from Table 6, our AdamX achieved the highest mIoU of $37.81\%$.
@@ 5.4 Ablation study results and analysis
Figure 4 visualizes the second-order moment estimate exponential decay rates in different optimization algorithms. The ablation experiments in this section focused on comparing the different second-order moment estimate exponential decay rates shown in this figure. Therefore, the $\overset{\frown}{\beta}_{2,t}$ in Adam was replaced with the $\overset{\frown}{\beta}_{2,t}$ from Adafactor [20], AdaX, and AdamX, respectively, to verify which second-order moment estimate exponential decay rate is superior.
@@
Optimal performance comparison. Table 7 shows the ablation study results of the second-order moment estimate exponential decay rates on the CIFAR-100 dataset with the benchmark model EfficientNetV2-M. Table 8 shows the ablation study results of the second-order moment estimate exponential decay rates on the CIFAR-100 dataset with the benchmark model SwinV2-S. Table 9 shows the ablation study results of the second-order moment estimate exponential decay rates on the VOC2007 test dataset with the benchmark model YOLOv7-tiny. Table 10 shows the ablation study results of the second-order moment estimate exponential decay rates on the Semantic Boundaries dataset with the benchmark model U-Net. Based on the above tables, the $\overset{\frown}{\beta}_{2,t}$ of AdamX is statistically the best, further verifying the two excellent properties:
@@ 6 Comparison results and analysis of optimization algorithms of different type
Then, by comparing our AdamX with different types of optimization algorithms (momentum SGD [8], RAdam [21], Lion [10], SophiaG [11]), the superiority of AdamX in performance was more comprehensively verified. To ensure the fairest comparison possible, all optimization algorithms followed their respective optimal hyperparameter configurations as much as possible. The total number of training epochs was set to $80$, and the batch size was set to $64$.
Table 11 shows the comparative experimental results on the benchmark model EfficientNetV2-L. For momentum SGD, the momentum was set to $0.9$, the peak learning rate was set to $0.1$, the minimum learning rate was set to $1\times 10^{-3}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $1\times 10^{-3}$; for Lion, the peak learning rate was set to $1\times 10^{-3}$, the minimum learning rate was set to $1\times 10^{-5}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.1$; for RAdam, SophiaG, and AdamX, the peak learning rate was set to $0.01$, the minimum learning rate was set to $1\times 10^{-4}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.01$. As seen in Table 11, our AdamX achieved the lowest top-1 error rate of $35.66\%$ and the lowest top-5 error rate of $12.37\%$ on EfficientNetV2-L.
@@
AdamX (Ours) 35.66 12.37
Table 12 shows the comparative experimental results with the baseline model being ConvNeXt-B. To adapt to the image resolution in CIFAR, ConvNeXt-B only retained the last three downsamplings, all convolutional kernel sizes were set to $(3,3)$, and the padding was set to $(1,1)$. For momentum SGD, the momentum was set to $0.9$, the peak learning rate was set to $0.01$, the minimum learning rate was set to $1\times 10^{-4}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $1\times 10^{-3}$; for Lion, the peak learning rate was set to $1\times 10^{-4}$, the minimum learning rate was set to $1\times 10^{-6}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.1$; for RAdam, SophiaG, and AdamX, the peak learning rate was set to $1\times 10^{-3}$, the minimum learning rate was set to $1\times 10^{-5}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.01$. As shown in Table 11, our AdamX achieved the lowest top-1 error rate of $32.87\%$ and the lowest top-5 error rate of $10.72\%$ on ConvNeXt-B.
@@ 7 Conclusion
This paper proposes a novel AdamX, whose core innovation lies in the novel $\overset{\frown}{\beta}_{2,t}=\left(1-\beta_{2}^{(1-\beta_{2})(t-1)}\right)/\left(1-\beta_{2}^{(1-\beta_{2})t}\right)$, which gradually reduces the correction strength of the learning step as the training progresses, thus degenerating to momentum SGD during the training plateau phase. The experimental results show that the proposed $\overset{\frown}{\beta}_{2,t}$ is superior to the $\overset{\frown}{\beta}_{2,t}$ of the existing optimization algorithms, and AdamX also performs better than the existing optimization algorithms. Future work includes whether the computational requirements of AdamX can be reduced.

### v3

# AdamNX: An Adam improvement algorithm based on a novel exponential decay mechanism for the second-order moment estimate

@@ Abstract
Since the 21st century, artificial intelligence has been leading a new round of industrial revolution. Under the training framework, the optimization algorithm aims to stably converge high-dimensional optimization to local and even global minima. Entering the era of large language models, although the scale of model parameters and data has increased, Adam remains the mainstream optimization algorithm. However, compared with stochastic gradient descent (SGD) based optimization algorithms, Adam is more likely to converge to non-flat minima. To address this issue, the AdamNX algorithm is proposed. Its core innovation lies in the proposition of a novel type of second-order moment estimation exponential decay rate, which gradually weakens the learning step correction strength as training progresses, and degrades to momentum SGD in the stable training period, thereby improving the stability of training in the stable period and possibly enhancing generalization ability. Experimental results show that our second-order moment estimation exponential decay rate is better than the current second-order moment estimation exponential decay rate, and AdamNX can stably outperform Adam and its variants in terms of performance. Our code is open-sourced at https://github.com/mengzhu0308/AdamNX.
@@
Optimization algorithms , Adam , SGD , AdamNX , Exponential decay rate of second-order moment estimate.
@@ 1 Introduction
With the advent of the era of large language models (LLM), the currently mainstream optimization algorithm remains Adam. Therefore, it still holds research value to make some minor modifications to Adam. The main criticisms of Adam are twofold: (1) it requires additional sets of parameters to cache the first-order and second-order moments. (2) It tends to converge to sharp minima, which have poorer generalization capabilities [13, 14]. In response to the issue of Adam’s slightly inferior generalization, this paper proposes the AdamNX algorithm (see Algorithm 1), whose core innovation lies in the proposition of a novel exponential decay rate for the second-order moment estimation, which controls the gradual weakening of the learning step correction strength as training progresses, thereby degrading to momentum SGD in the stable training period.
The remainder of this paper is organized as follows. Section 2 reviews the existing research on optimization algorithms in deep learning and introduces the motivation of this study. Section 3 provides a detailed analysis of the problem and proposes the corresponding solution. Sections 4, 5, and 6 jointly verify the superiority of AdamNX through experiments. The final section summarizes the content of the entire paper.
@@ 3 Proposed algorithm
This section first analyzes the impact of $\beta_{2}$ on the variance of the second-order moment estimation in Adam, semi-qualitatively and semi-quantitatively proving that when $t\to\infty,\beta_{2}\to 1$, Adam can degrade into momentum SGD, thereby reducing the “jitter” phenomenon in the stable training period. Then, to address the contradiction in Adam that $\beta_{2}$ cannot be too large or too small, a novel exponential decay rate for the second-order moment estimation is proposed, along with the corresponding AdamNX optimization algorithm. Finally, AdamNX and Adam are compared once again.
@@ 3.1 Impact of $\beta_{2}$ on the variance of the second-order moment estimation in Adam
3.2 Novel exponential decay rate for the second-order moment estimation and AdamNX
@@
and correspondingly proposes the AdamNX algorithm, as shown in Algorithm 1. To align with Adam’s default settings: $\beta_{1}=0.9$, $\beta_{2}=0.999$, AdamNX is set to: $\beta_{1}=0.9$, $\beta_{2}=0.99$. This allows hyperparameters tuned for Adam to be better transferred to AdamNX. Figure 1 visualizes $\overset{\frown}{\beta}_{1,t}$ and $\overset{\frown}{\beta}_{2,t}$ in AdamNX.
@@ 13    $\mathcal{L}_{0}\leftarrow\mathcal{L}_{t}$;
Algorithm 1 AdamNX algorithm
Figure 1: Functional curves of $\overset{\frown}{\beta}_{1,t}$ and $\overset{\frown}{\beta}_{2,t}$ versus iteration number $t$ in AdamNX
3.3 AdamNX v.s. Adam
@@
Comparing Equations (12) and (14), the core difference is that when $t\to\infty$, AdamNX’s $\overset{\frown}{\beta}_{2,t}\to 1$, thus AdamNX degenerates to momentum SGD, guaranteeing more stability in the later training stage. Figure 2 further visualizes the comparison between AdamNX’s $\overset{\frown}{\beta}_{2,t}$ and Adam’s $\overset{\frown}{\beta}_{2,t}$ for different $\beta_{2}$ values.
Figure 2: Comparison between AdamNX’s $\overset{\frown}{\beta}_{2,t}$ and Adam’s $\overset{\frown}{\beta}_{2,t}$
@@ 5 Comparison results and analysis of optimization algorithms of the same type
Firstly, the superiority of AdamNX in performance was verified by comparing it with optimization algorithms of the same type (Adam [2] and AdaX [22]) (These optimization algorithms can use the same training configuration). Experiments were conducted on the CIFAR-100 [24], PASCAL VOC [25], and Semantic Boundaries [26] datasets, with EfficientNetV2 [27] and SwinV2 [28] selected as the benchmark models for image classification, YOLOv7-tiny [29] as the benchmark model for object detection, and U-Net [30] as the benchmark model for semantic segmentation.
@@ 5.1 Image classification experimental results and analysis
Lowest error rate comparison. Table 3 shows the comparative experimental results on the benchmark model EfficientNetV2-M. The total number of training epochs was set to $80$, the batch size was set to $64$, the peak learning rate was set to $3\times 10^{-3}$, the minimum learning rate was set to $3\times 10^{-5}$, and the training $t_{1}$ was set to $37,488$. As seen in Table 3, our AdamNX achieved the lowest top-1 error rate of $34.27\%$ and the lowest top-5 error rate of $11.40\%$ on EfficientNetV2-M. Table 4 shows the comparative experimental results on the benchmark model SwinV2-S. The total number of training epochs was set to $80$, the batch size was set to $64$, and the fixed learning rate was set to $2.2\times 10^{-5}$. As seen in Table 4, our AdamNX also achieved the lowest top-1 error rate of $39.42\%$ and the lowest top-5 error rate of $14.68\%$ on SwinV2-S.
@@ 5.2 Object detection experimental results and analysis
Highest mAP comparison. Table 5 shows the comparative experimental results on the VOC2007 test benchmark dataset. The total number of training epochs was set to $80$, the batch size was set to $32$, the peak learning rate was set to $3\times 10^{-3}$, the minimum learning rate was set to $3\times 10^{-5}$, and the training $t_{1}$ was set to $25,680$. As seen in Table 5, our AdamNX achieved the highest mAP@0.5 of $52.18\%$ and the highest mAP@0.75 of $28.21\%$.
@@ 5.3 Semantic segmentation experimental results and analysis
Highest mIoU comparison. Table 6 shows the comparative experimental results on the Semantic Boundaries benchmark dataset. The total number of training epochs was set to $80$, the batch size was set to $32$, the peak learning rate was set to $1\times 10^{-3}$, the minimum learning rate was set to $1\times 10^{-5}$, and the training $t_{1}$ was set to $12,720$. As can be seen from Table 6, our AdamNX achieved the highest mIoU of $37.81\%$.
@@ 5.4 Ablation study results and analysis
Figure 4 visualizes the second-order moment estimate exponential decay rates in different optimization algorithms. The ablation experiments in this section focused on comparing the different second-order moment estimate exponential decay rates shown in this figure. Therefore, the $\overset{\frown}{\beta}_{2,t}$ in Adam was replaced with the $\overset{\frown}{\beta}_{2,t}$ from Adafactor [20], AdaX, and AdamNX, respectively, to verify which second-order moment estimate exponential decay rate is superior.
@@
Optimal performance comparison. Table 7 shows the ablation study results of the second-order moment estimate exponential decay rates on the CIFAR-100 dataset with the benchmark model EfficientNetV2-M. Table 8 shows the ablation study results of the second-order moment estimate exponential decay rates on the CIFAR-100 dataset with the benchmark model SwinV2-S. Table 9 shows the ablation study results of the second-order moment estimate exponential decay rates on the VOC2007 test dataset with the benchmark model YOLOv7-tiny. Table 10 shows the ablation study results of the second-order moment estimate exponential decay rates on the Semantic Boundaries dataset with the benchmark model U-Net. Based on the above tables, the $\overset{\frown}{\beta}_{2,t}$ of AdamNX is statistically the best, further verifying the two excellent properties:
@@ 6 Comparison results and analysis of optimization algorithms of different type
Then, by comparing our AdamNX with different types of optimization algorithms (momentum SGD [8], RAdam [21], Lion [10], SophiaG [11]), the superiority of AdamNX in performance was more comprehensively verified. To ensure the fairest comparison possible, all optimization algorithms followed their respective optimal hyperparameter configurations as much as possible. The total number of training epochs was set to $80$, and the batch size was set to $64$.
Table 11 shows the comparative experimental results on the benchmark model EfficientNetV2-L. For momentum SGD, the momentum was set to $0.9$, the peak learning rate was set to $0.1$, the minimum learning rate was set to $1\times 10^{-3}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $1\times 10^{-3}$; for Lion, the peak learning rate was set to $1\times 10^{-3}$, the minimum learning rate was set to $1\times 10^{-5}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.1$; for RAdam, SophiaG, and AdamNX, the peak learning rate was set to $0.01$, the minimum learning rate was set to $1\times 10^{-4}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.01$. As seen in Table 11, our AdamNX achieved the lowest top-1 error rate of $35.66\%$ and the lowest top-5 error rate of $12.37\%$ on EfficientNetV2-L.
@@
AdamNX (Ours) 35.66 12.37
Table 12 shows the comparative experimental results with the baseline model being ConvNeXt-B. To adapt to the image resolution in CIFAR, ConvNeXt-B only retained the last three downsamplings, all convolutional kernel sizes were set to $(3,3)$, and the padding was set to $(1,1)$. For momentum SGD, the momentum was set to $0.9$, the peak learning rate was set to $0.01$, the minimum learning rate was set to $1\times 10^{-4}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $1\times 10^{-3}$; for Lion, the peak learning rate was set to $1\times 10^{-4}$, the minimum learning rate was set to $1\times 10^{-6}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.1$; for RAdam, SophiaG, and AdamNX, the peak learning rate was set to $1\times 10^{-3}$, the minimum learning rate was set to $1\times 10^{-5}$, $t_{1}$ was set to $37,488$, and the weight decay rate was set to $0.01$. As shown in Table 11, our AdamNX achieved the lowest top-1 error rate of $32.87\%$ and the lowest top-5 error rate of $10.72\%$ on ConvNeXt-B.
@@ 7 Conclusion
This paper proposes a novel AdamNX, whose core innovation lies in the novel $\overset{\frown}{\beta}_{2,t}=\left(1-\beta_{2}^{(1-\beta_{2})(t-1)}\right)/\left(1-\beta_{2}^{(1-\beta_{2})t}\right)$, which gradually reduces the correction strength of the learning step as the training progresses, thus degenerating to momentum SGD during the training plateau phase. The experimental results show that the proposed $\overset{\frown}{\beta}_{2,t}$ is superior to the $\overset{\frown}{\beta}_{2,t}$ of the existing optimization algorithms, and AdamNX also performs better than the existing optimization algorithms. Future work includes whether the computational requirements of AdamNX can be reduced.